How many liters of a 50% acid solution must be mixed with a 30% acid solution to get 180L of a 40% acid solution?
Answer by mananth(16946)


percent ---------------- quantity
Acid type I 50.00% ---------------- x liters
acid type II 30.00% ------ 180 - x liters
Mixture 40.00% ---------------- 180
180
50.00% x + 30.00% ( 180 - x ) = 40.00% * 180
50 x + 30 ( 180 - x ) = 7200
50 x + 5400 - 30 x = 7200
50 x - 30 x = 7200 - -5400
20 x = 1800
/ 20
x = 90 liters 50.00% Acid type I
90 liters 30.00% acid type II

m.ananth@hotmail.ca